Answer the following questions regarding heats of fomation and combustion:

The balanced equation for the combustion of salicylic acid is given to be:

C7H6O3(s) + 7O2(g) --> 7CO2(g) + 3H2O(l)

A. Calculate the heat of formation (\Delta H_f) of salicylic acid with the following data provided:

For H2O(l): \Delta H_f = -286.2 kJ mol^-^1

For CO2(g): \Delta H_f = -393.3 kJ mol^-^1

B. Calculate the heat of formation of benzoic acid using the data for its heat of combustion obtained from the literature. Assume 1% accuracy for any thermochemical data used.

C. A power station is being designed to produce 1200 MW of electricity using the coal supplied in a particular experiment. If the design efficiency of the plant is 30%, calculate the daily consumption of coal and the daily mass of ash for disposal. Given that the coal used in the experiment has an actual mass of 0.4937 g and an ash content after bomb calorimeter of 97.3 %. The heat of combustion from the bomb calorimeter of coal was found to be 22.45 MJ/kg and the heat released was calculated to be -22.10 kJ/g.

D. Calorimeters other than a bomb exist for determining thermodynamic parameters of a chemical reaction. Identify one of these and outline how and what parameters might be measured.
Answer #1


A)

C7H6O3(s) + 7O2(g) --> 7CO2(g) + 3H2O(l) DHrxn = -3020 kJ/mol

DHrxn = (7*DHf0CO2+3*DHfOH2O)-(7*DHfOO2+1*DHfOc7h6o3)

-3020 = (7*-393.3+3*-286.2)-(7*0+x)

X = DHf0 = enthalpy of formtaion of salisilic acid(C7H6O3) = -591.7 kj/mol


B)

2C6H5COOH(s) + 15O2(g) --> 14CO2(g) + 6H2O(l) DHrxn = -3226.6 kj/mol

DHrxn = (14*DHf0CO2+6*DHfOH2O)-(15*DHfOO2+1*2*DHfOC6H5COOH)

DHrxn = (14*DHf0CO2+6*DHfOH2O)-(15*DHfOO2+2*DHfOC6H5COOH)

2*-3226.6 = (14*-393.3+6*-286.2)-(15*0+2*x)

x = DHf0 = enthalpy of formtaion of benzoicacid = -385.1 kj/mol

X = DHf0 = enthalpy of formtaion of C7H6O3 = -591.7 kj/mol